Acetone, C3H6, is the main ingredient of nail polish remover. A solution is made up




by adding 35.OmL of acetone (d=0.790g/mL) to 50.0mL of ethyl alcohol, C2H60




(d=0.789g/mL). Assuming volumes are additive, calculate (a) the mass percent of




acetone in the solution. (b) the volume percent of ethyl alcohol in the solution. (c) the




mole fraction of acetone in the solution.

1
Expert's answer
2022-04-06T09:31:04-0400

density = m / V

m (C3H6) = 35 x 0.79 = 27.7 g

m (C2H6O) = 50 x 0.789 = 39.5 g

a) Mass percent acetone in the solution = (100 x 27.7) / (27.7 + 39.5) = 41%


b) the volume percent of ethyl alcohol in the solution: (100 x 50) / (35 + 50) = 59%


(c) n = m / M

M (C3H6) = 42.1 g/mol

M (C2H6O) = 46.1 g/mol

n (C3H6) = 27.7 / 42.1 = 0.66 mol

n (C2H6O) = 39.5 / 46.1 = 0.87 mol

The mole fraction of acetone in the solution: 0.66 / (0.66 + 0.87) = 0.43